OpenStudy (anonymous):

Which value of x is a solution on 5/x= x+13/6? a. -2 c. -10 b. -3 d. -15

OpenStudy (shubhamsrg):

ever tried "hit and trial" ?

OpenStudy (anonymous):

i dont know i just made an account right now

OpenStudy (shubhamsrg):

...

jimthompson5910 (jim_thompson5910):

5/x = (x+13)/6 5*6 = x(x+13) 30 = x^2 + 13x 0 = x^2 + 13x - 30 x^2 + 13x - 30 = 0 (x - 2)(x + 15) = 0 Do you see how to get the two solutions?

OpenStudy (anonymous):

Ohhh i do! thanks i so get it now :)

jimthompson5910 (jim_thompson5910):

yw